# Text-Speech Language Models with Improved Cross-Modal Transfer by Aligning Abstraction Levels

### v1

@@ 4.2 Evaluation
We also report text performance on MMLU (Hendrycks et al., 2021) before (pre) and after (post) speech training to assess whether fine-tuning causes forgetting of text capabilities, as seen in other TSLMs (Nguyen et al., 2024; Défossez et al., 2024). We evaluate MMLU following the guidelines for the SmolLM models: https://huggingface.co/HuggingFaceFW/ablation-model-fineweb-edu#evaluation.
Baselines. We compare SmolTolk to models trained with the same text LM backbones and data but using regular vocabulary expansion, referring to these as baselines. We also compare against state-of-the-art TSLMs: Spirit LM (Nguyen et al., 2024), Moshi (Défossez et al., 2024), and the 1.5-billion and 9-billion models from Zeng et al. (2024).
@@
Zeng et al. (2024) 1.5B 1.5B 1T 77.5 81.4 90.1 55.4 58.6 64.0
Zeng et al. (2024) 9B 9B 1T 83.0 85.0 93.6 62.4 63.2 76.3
@@
4.3 Data
Speech datasets. We use a collection of publicly available English speech datasets for training: LibriSpeech (Panayotov et al., 2015), LibriLight (Kahn et al., 2020), SWC (Baumann et al., 2019), Tedlium (Hernandez et al., 2018), People’s Speech (Galvez et al., 2021), Vox Populi (Wang et al., 2021), and sTinyStories (Cuervo & Marxer, 2024). These datasets contain a total of 10.89 billion speech tokens.
Text datasets. We use a 12-billion-token subset of the SmolLM corpus (Ben Allal et al., 2024). Unlike Nguyen et al. (2024), we include math and code data, aiming to better preserve text capabilities. Our data distribution matches that used for pre-training SmolLM models, as reported in https://github.com/huggingface/smollm/blob/main/pre-training/.
Text-Speech datasets. We use the forced aligner from Pratap et al. (2024) to obtain word alignments for the LibriHeavy (Kang et al., 2024), sTinyStories, and SWC datasets. Interleaved samples are generated on the fly during batch sampling by randomly switching modalities within the input sequence. Following Nguyen et al. (2024), we randomly select word spans so that each text sequence contains 10–30 words and each speech sequence 5–15 words, balancing the proportion of speech and text tokens in each sample.
For more details on the datasets see Appendix B.
5 Experiments and results
@@
Table 1 presents the benchmark results. For brevity, we report only the results for the 1.7-billion baseline. Smaller baseline models underperformed relative to the larger one. An expanded results table is provided in Appendix D.
Our method significantly outperforms the baseline using regular vocabulary expansion across all tasks. One might attribute this difference to the parameter increase induced by the added modules. To better account for this factor, Figure 2 shows the scaling behavior of the negative log-likelihood (NLL) on the LibriSpeech dev set and the tStoryCloze benchmark as a function of compute, which helps adjust for model size differences and enables a fairer comparison. The figure demonstrates that, across the entire compute range, our models consistently outperform their respective baselines. Notably, SmolTolk-150M surpasses the 360-million baseline and rivals or surpasses the 1.7-billion baseline, suggesting that factors beyond model size drive the performance difference. Interestingly, in the baselines, cross-modal alignment scales only in the 1.7-billion parameter model. In contrast, our architecture demonstrates this capability even at smaller model sizes, indicating a more efficient approach to cross-modal learning. Scaling behavior for other metrics is presented in Appendix D.
@@
Table 2: Ablation Study. “–” denotes removal. ”–Dyn. pooling” uses fixed learned weights instead of dynamic ones from the layer selector, while ”–Layer pooling” entirely disables multi-layer pooling, relying only on the last text LM layer.
[…]
Our setup, including the baselines, exhibits less deterioration in text MMLU performance relative to other TSLMs. We attribute this to our decision to use a text fine-tuning distribution that matches the one used during pre-training.
5.1 Ablation study
Overall, the results indicate that our design choices enhance multimodal performance. To better understand each component’s contribution, we conduct an ablation study in Table 2 on the medium-size model by systematically removing elements and evaluating their impact.
We observe that removing any component degrades performance across most metrics, confirming the importance of our design choices. Eliminating all adapters results in the steepest drop—especially in cross-modal transfer, highlighting their role in bridging representations. The input adapter seems to be of greater importance for cross-modal transfer than the output adapter, underscoring the importance of merging modalities early in processing. Layer pooling also provide consistent gains, demonstrating the benefits of allowing the model to use multiple abstraction levels for speech language modeling. Removing the residual connection also causes a consistent decrease in performance.
Figure 3: Intrinsic dimensionality of representations across layers.
Figure 4: Fraction of variance explained by projecting one modality onto the other’s top 50 principal components across layers.
5.2 Representation analysis
To gain deeper insight into our design choices, we analyze learned representations across model variants. We investigate how the introduced architectural components influence feature abstraction and alignment between text and speech. We focus on two aspects: (1) the model’s capacity to abstract high-level features and (2) the shared structure of text and speech representations.
@@ 5.2 Representation analysis
To assess (1), we follow Valeriani et al. (2023) and use the intrinsic dimensionality of representations as a proxy for compositionality. To evaluate (2), we compute the principal components of speech and text representations on paired data and measure how much variance is explained when projecting one modality’s representations onto the other’s principal components. The intuition is that knowledge transfer can be quantified by how much the model utilizes the same subspaces to represent equivalent data across modalities. We apply these analyses to the medium size models, including the architectural ablations from Table 2. For details on the experimental setup, see Appendix C. The results are shown in Figures 4 and 4.
From Figure 4, we observe that different components of our architecture significantly impact the model’s ability to compose higher-abstraction features. All our models, except the one without layer pooling, achieve a higher intrinsic dimension, suggesting that layer pooling is essential for enabling compositionality. The effect of the adapters is also evident: the absence of input adapters leads to lower compositionality in earlier layers, while the absence of output adapters results in an overall reduction in intrinsic dimension.
Figure 4 supports similar conclusions, showing that our architecture enables the highest degree of subspace overlap between modalities. As before, adapters and layer pooling are crucial for cross-modal transfer. The absence of input adapters leads to low subspace overlap in earlier layers, while the absence of output adapters reduces overlap in later layers. We hypothesize that this occurs because the model repurposes later layers to produce representations predictive of upcoming speech tokens. In this study the speech input residual has minimal effect.
@@
What is dynamic layer pooling learning?. Figure 5 shows the weights assigned by the layer selector $S$ across layers (top) and for the last layer—specialized in next token prediction—(bottom) across a speech input, alongside word boundaries. The layer weights follow a pattern consistent with our hypothesis, switching between low-level representations and those predictive of next words. Notably, spikes in last-layer weights often align with word boundaries, suggesting the model leverages next-word predictive representations at these points. To validate this, we compute a speech word segmentation score using a peak detector on the last layer’s attention weights as a word boundary predictor (see Appendix C for details). Table 3 compares our approach with SCPC (Bhati et al., 2022), a state-of-the-art unsupervised speech segmentation model, on the TIMIT test split (Garofolo et al., 1993). Our method outperforms SCPC, supporting our hypothesis on dynamic pooling behavior. However, we do not claim our segmentation method is unsupervised, as the model is trained on interleaved data that implicitly provides word boundary information. Interestingly, segmentation performance scales inversely with model size. We hypothesize that this may be because the greater capacity of output adapters in larger models allows them to rely less on backbone features. For segmentation examples of other sequences and models see Appendix D.
@@
Figure 5: Selector $S$ layer weights across a speech input sequence for SmolTolk-2B. Vertical bars indicate word boundaries.
@@
Table 3: Word segmentation scores on TIMIT test set using dynamic pooling last layer weights as word boundary predictors.
@@ 6 Related work
Chou et al. (2023), Spirit LM (Nguyen et al., 2024), and Zeng et al. (2024) train TSLMs by jointly modeling text, speech, and interleaved text-speech without using text as an intermediate representation for speech generation. We adopt a similar training approach but propose a different speech modeling architecture that enhances multimodal performance. Moreover, compared to Spirit LM and Zeng et al. (2024), we use smaller, less diverse open speech datasets and far fewer computational resources while achieving comparable or, at times, superior performance.
Other state-of-the-art TSLMs follow a hierarchical process, first generating text as an intermediate representation to condition speech generation, akin to TTS. Notable works include SpeechGPT (Zhang et al., 2023), Moshi (Défossez et al., 2024), and Llama-Omni (Fang et al., 2024). Unlike these, our model does not rely on text generation before speech synthesis. Llama-Omni is particularly relevant, as it employs input speech adapters with temporal downsampling and an MLP layer, and an output adapter with temporal upsampling and transformer decoder layers. In contrast, our adapters avoid temporal downsampling/upsampling, eliminating latency, and use transformer decoder layers in the input adapter as well, as we have evidence from Valeriani et al. (2023); Cheng et al. (2025) that these layers can perform the composition processes we posit are necessary.
A closely related work architecturally is LAST (Turetzky & Adi, 2024), where adapters with the same architecture as ours feed and generate speech from a text LM, primarily to learn a speech tokenization from HuBERT representations. Unlike LAST, we apply multi-layer pooling before the output adapter, perform multimodal training, target cross-modal capabilities, and achieve state-of-the-art performance.
7 Discussion and future work
7.1 Implications
On speech language modeling. We believe our results make a strong case for future TSLM research to adopt our proposed architecture. Our method can complement other approaches to improving speech language modeling, such as large-scale diverse datasets (Zeng et al., 2024), additional losses (Lin et al., 2024), or coarser speech tokenization (Baade et al., 2025). The latter may better interface with the text LM and reduce the load—and thus the required capacity—of adapters.
Our dynamic multi-layer pooling method may also be relevant for textless speech LM research, as a similar representation hierarchy likely emerges in speech LMs, reflecting speech’s inherent feature hierarchy. Thus, speech predictive layers should adaptively switch between attending to low- and high-level representations to facilitate high-level feature abstraction, as suggested by the experiments in Section 5.2.
On interfacing other modalities to text LMs. Overall, our results support our initial hypotheses, highlighting the importance of accounting for feature compositionality when adapting text LMs for speech modeling. We believe this insight may extend to other modalities, as multimodal generative LMs use similar vocabulary expansion strategies, e.g., in image modeling (Wang et al., 2022; Aghajanyan et al., 2022; Chameleon Team, 2024). Although models like Chameleon (Chameleon Team, 2024) tout the absence of modality-specific modules as a strength, our findings suggest this may come at the cost of cross-modal alignment.
More broadly, we think of text as the most compressed representation humans use to interact with the world, closest to a Platonic representation of abstract concepts, following the thought-provoking idea from Huh et al. (2024). As such, we believe most other perceptual modalities naturally exist at lower abstraction levels (i.e., less compressed) and require composition at the input and refinement at the output to properly interface with text LMs. Therefore, we suggest that applying the ideas used in this work for speech to other modalities is a worthy direction for future research.
7.2 Limitations
@@
Direct applications. Since our model is not instruction fine-tuned, its applicability to common speech LM tasks, such as spoken conversational agents, is limited. Future work could explore instruction tuning and broader evaluations. Additionally, using linguistic tokens makes our model less expressive—one of speech LMs’ key advantages over cascaded approaches combining text LMs with ASR and TTS. Future work could apply our architecture to model more expressive speech representations.
Scale. Our models are small by current standards, and the observed phenomena may differ at larger scale. Figure 2 shows that cross-modal transfer scaling tends to improve in larger models without our architecture. This suggests that simpler models could match our performance at scale, echoing The Bitter Lesson (Sutton, 2019). However, in line with The Bitter Lesson, our design imposes no additional priors, allowing the model to learn to effectively revert to the simpler baseline architecture if optimal. Therefore, we do not expect it to bottleneck performance. Instead, given the observed gains, it is likely to remain beneficial.
Novelty. While perhaps not explicitly stated, the hypotheses studied in this work clearly underlie many prior multimodal LM studies using adapters, as they are fairly intuitive and perhaps even obvious. However, previous works doing vocabulary expansion gave little consideration to these aspects, and we show this may have significantly harmed performance. Moreover, our ablation and representation analyses strengthen the understanding of the underlying phenomena. Additionally, to the best of our knowledge, we are the first to argue for and demonstrate the usefulness of multi-layer representations in speech language modeling.
@@ 8 Conclusions
We introduced a more effective approach to text-speech language modeling by enhancing vocabulary expansion with speech-specific adapters and dynamic layer pooling. These components improve abstraction alignment across model layers, enabling more effective cross-modal transfer. Our evaluations and representation analyses show that this method better integrates speech into text LMs.
Our SmolTolk models achieve state-of-the-art performance using far fewer computational resources and smaller, less diverse datasets than previous works. Notably, SmolTolk-2B rivals or surpasses much larger TSLMs. These results underscore the importance of hierarchical feature compositionality in multimodal learning. Beyond speech, our findings suggest that aligning feature abstraction levels may be key to adapting text LMs for other modalities.
@@
Santiago proposed and led the project, defining hypotheses, model and experimental design, and developed most of the training, model and analysis codebase. Adel played a key role in data preprocessing and synthesis, text and cross-modal evaluations, and model inference, while also contributing insights on model design. He maintained the codebase with Santiago’s assistance. Ricard provided supervision throughout and contributed significantly to the data pipeline, experimental codebase, and analysis. Yanis assisted with experiment implementation, execution, and discussions. Other authors contributed to early-stage discussions and supervision during the JSALT 2024 workshop.
@@ Appendix A Models details
Table 4 describes the SmolLM (Ben Allal et al., 2024) backbones architectural hyperparameters. All SmolTolk models use two layer input and output adapters with the same architecture as the backbone layers. Table 5 describes the resulting models after vocabulary expansion.
@@
As mentioned in Section 3.1, we use a linear layer with bias as the dynamic layer selector $S$. We experimented with simple non-linear MLP selectors; however, these were prone to collapse and resulted in worse overall performance. That said, a more carefully designed non-linear selector could potentially perform better. We also explored alternative ways to define the contextual representation $\mathbf{c}_{i}^{\prime}$ (Equation 3) used as input for the layer selector. Instead of a learned weighted average, we tried concatenating low-dimensional linear projections from each layer’s representations, but this performed worse.
@@
For our experiments, we used NVIDIA H100 GPU nodes, each featuring four NVIDIA H100 80GB SXM5 GPUs, dual Intel Sapphire Rapids 48-core processors, 512GB of RAM, and four NVIDIA ConnectX-7 400Gb/s InfiniBand network adapters. SmolTolk-2B used a per-GPU batch size of 8 million tokens and was trained on 64 GPUs across 16 nodes, taking approximately 10 hours to process 32 billion tokens. All models were trained using bfloat16 mixed precision with FlashAttention-2 (Dao, 2024) and PyTorch compile.

### v2

@@ 4.2 Evaluation
We also report text performance on MMLU (Hendrycks et al., 2021) before (pre) and after (post) speech training to assess whether fine-tuning causes forgetting of text capabilities, as seen in other TSLMs (Nguyen et al., 2024; Défossez et al., 2024). MMLU is evaluated following the guidelines for the SmolLM models [^2] https://huggingface.co/HuggingFaceFW/ablation-model-fineweb-edu#evaluation.
Baselines. We compare SmolTolk to models trained with the same text LM backbones and data but using early fusion/fission, referring to these simply as baselines. We also compare against state-of-the-art early fusion/fission TSLMs: SpiritLM (Nguyen et al., 2024), Moshi (Défossez et al., 2024), and the 1.5-billion and 9-billion GLM-4-Voice models (Zeng et al., 2024).
4.3 Data
Speech datasets. We train on a collection of public English speech datasets: LibriSpeech (Panayotov et al., 2015), LibriLight (Kahn et al., 2020), SWC (Baumann et al., 2019), Tedlium (Hernandez et al., 2018), People (Galvez et al., 2021), Vox Populi (Wang et al., 2021), and sTinyStories (Cuervo and Marxer, 2024), totaling 160k hours of audio (10.89B speech tokens).
Text datasets. We use a 12-billion-token subset of the SmolLM corpus (Allal et al., 2025). Unlike Nguyen et al. (2024), we include math and code data, aiming to better preserve text capabilities. Our data distribution matches that used for pre-training SmolLM models [^3] As reported in https://github.com/huggingface/smollm/blob/main/pre-training/.
Text-Speech datasets. We use the forced aligner from Pratap et al. (2024) to obtain word-level time alignments for the LibriHeavy (Kang et al., 2024), sTinyStories, and SWC datasets. Interleaved samples are generated on the fly during batch sampling by randomly switching modalities within the input sequence. Following Nguyen et al. (2024), we randomly select word spans so that each text sequence contains 10–30 words and each speech sequence 5–15 words, balancing the proportion of speech and text tokens in each sample.
For more detail on the datasets see Appendix B.
5 Experiments and results
Table 1 presents the benchmark results. For brevity, we report only the results for the 1.7-billion baseline. Smaller baseline models underperformed relative to the larger one. An expanded results table is provided in Appendix D.5. See Appendix D.3 for cross-modal generation samples.
Our method significantly outperforms the early fusion/fission baselines across all tasks. One might attribute this difference to the parameter increase induced by the added modules. To better account for this factor, Figure 2 shows the scaling behavior of the negative log-likelihood (NLL) on the LibriSpeech dev set and the tStoryCloze benchmark as a function of compute, which helps adjust for model size differences and enables a fairer comparison. The figure demonstrates that, across the entire compute range, our models consistently outperform their respective baselines. Notably, SmolTolk-150M surpasses the 360-million baseline and rivals or surpasses the 1.7-billion baseline, proving that factors beyond model size drive the performance difference. Interestingly, in the baselines, cross-modal alignment scales only in the 1.7-billion parameter model. In contrast, our architecture demonstrates this capability even at smaller model sizes, indicating a more efficient approach to cross-modal learning. Scaling behavior for other metrics is presented in Appendix D.6.
@@ 5 Experiments and results
GLM-4-Voice (Zeng et al., 2024) 1.5B 1.5B 1T 77.5 81.4 90.1 55.4 58.6 64.0
GLM-4-Voice (Zeng et al., 2024) 9B 9B 1T 83.0 85.0 93.6 62.4 63.2 76.3
@@
Table 1: Downstream evaluations. The best model in each task is shown in bold and underlined. The second best is shown in bold. For Spirit LM we report the results for the open-weights version. For other models we present the results reported by the authors. For MMLU, the best model is defined as the one exhibiting the least degradation.
@@
[…]
Our setup, including the baselines, exhibits less deterioration in text MMLU. We attribute this to our decision to use a text fine-tuning distribution that matches the one used during pre-training.
5.1 Ablation study
Overall, the results indicate that our design choices enhance multimodal performance. To better understand each component contribution, we conduct an ablation study in Table 2 on the medium-size model by systematically removing elements and evaluating their impact. We observe that removing any component degrades performance across most metrics, confirming the importance of our design choices. Eliminating all adapters results in the steepest drop—especially in cross-modal transfer, highlighting the importance of accounting for the compositionality hierarchy. The input adapter seems to be of greater importance for cross-modal transfer than the output adapter, underscoring the importance of late fusion. Multi-level representations and the residual connection also provide consistent gains, demonstrating the benefits of multi-level fission.
@@ 5.1 Ablation study
Table 2: Architecture ablation study. “–” denotes removal. "–Dyn. pooling" uses fixed learned weights instead of dynamic ones from the layer selector, while "–Layer pooling" entirely disables multi-layer pooling, relying only on the last text LM layer.
Figure 3: Ablations. Top: Intrinsic dimensionality of representations. Bottom: Fraction of variance explained by cross-modal projections.
5.2 Representation analyses
@@
The results above show that our proposed architecture improves multimodal performance but offer limited insight into the representation-level hypotheses (a), (b), and (c) from Section 1. To probe the mechanisms underlying these gains, we analyze learned representations across ablated model variants, focusing on two aspects: (1) the model’s ability to abstract high-level features, and (2) the alignment of text and speech representation spaces. To assess (1), we follow Valeriani et al. (2023) and use intrinsic dimensionality (ID) as a proxy for compositionality and semantic abstraction. To evaluate (2), we compute the top-k principal components of speech and text representations on paired data and measure how much variance is explained when projecting one onto the other. The intuition is that cross-modal transfer is reflected in the degree to which shared subspaces are used to represent semantically equivalent content. We apply these analyses to the medium-sized models and their ablations from Table 2. Experimental details are provided in Appendix C.
Figure 3 (top) shows that different components affect the model’s ability to compose higher-level features. The early fusion/fission baseline yields the lowest intrinsic dimensionality (ID). All models—except the one without multi-level representations—achieve higher ID, indicating that layer pooling is crucial for compositionality. This could be predicted from hypothesis (c): both low- and high-level features are needed for speech generation. Without multi-level representations, the model relies solely on the final layer, which must retain low-level information (see Appendix D.2 for evidence)—relevant for most steps—potentially hindering higher-level abstraction. Late fusion/fission also plays a role: removing input adapters lowers early-layer ID, while removing output adapters reduces ID overall.
Figure 3 (bottom) shows that multi-level fission and late fusion/fission also affect the subspace overlap between modalities. Notably, as with ID, the absence of multi-level representations yields the lowest degree of alignment (excluding the baseline). While we cannot claim a causal link between ID and alignment, the evidence is consistent with hypothesis (a): insufficient compositionality may prevent alignment. The absence of input adapters results in low overlap in earlier layers, supporting the idea that late fusion performs the required composition. Similarly, removing output adapters reduces overlap in later layers, which aligns with hypothesis (b): the model would repurpose later layers—originally optimized for next-word prediction—to generate representations for upcoming speech tokens. In this study, the speech input residual has minimal effect.
@@
What is dynamic layer pooling learning?. Hypothesis (c) yields a testable prediction: the model should alternate attention between low-level and next-word predictive representations, with the latter gaining weight near word boundaries. Figure 4 shows the weights assigned by the layer selector $S$ across layers (top) and for the final layer—specialized in next-token prediction—(bottom), over a speech input with word boundaries. Visual inspection suggests a pattern consistent with our hypothesis: the model shifts attention between low- and high-level features, with spikes in final-layer weights often aligning with word boundaries. To quantify this, we apply a peak detector to the final-layer weights and treat peaks as predicted word boundaries (see Appendix C for details). Table 3 compares this segmentation method to SCPC (Bhati et al., 2022), a state-of-the-art unsupervised speech segmentation model, on the TIMIT test split (Garofolo et al., 1993). Our approach outperforms SCPC, supporting hypothesis (c). Note that we do not claim our segmentation method is unsupervised, as training on interleaved text-speech data implicitly provides word boundary cues. Interestingly, segmentation quality decreases with model size—we hypothesize that larger models rely more on output adapters and less on backbone features. Additional examples are given in Appendix D.7.
@@
Figure 4: Selector $S$ layer weights across a speech input sequence for SmolTolk-2B. Vertical bars indicate word endings.
@@
Table 3: Word‐segmentation scores on the TIMIT test set using dynamic‐pooling last‐layer weights as word‐boundary predictors.
@@ 6 Related work
Several prior works adopt late fusion architectures to improve modality alignment. In the speech domain, relevant examples include Yu et al. (2024b); Chu et al. (2023); Tang et al. (2024). A more recent wave of TSLMs has focused on decoupling speech generation from the text LM backbone via additional speech-specific output modules (Xu et al., 2025) or modality-specific forward passes (Xie and Wu, 2024)—an idea aligned with our late fission approach. Unlike these models, we use simpler causal adapters that support a unified input-output representation (discrete speech tokens) naturally suited to auto-regressive language modeling, and we do not rely on paired text conditioning for speech generation. In addition, we introduce multi-level fission—a novel contribution to the best of our knowledge—and provide new insights from representation analyses into the role and impact of modality fusion and fission mechanisms.
In terms of training and evaluation methodology, our work is most closely aligned with Nguyen et al. (2024) and Zeng et al. (2024), with whom we compare. Compared to them, we use smaller, less diverse open speech datasets and significantly less computational resources, while achieving comparable—or in some cases, superior—performance.
7 Limitations
Scale. Our models are small by current standards, and the observed phenomena may differ at larger scales. Figure 2 shows that cross-modal transfer improves in larger models with early fusion/fission—suggesting that scale might help close the gap with our architecture. However, as shown in Appendix D.1 (Figure 5), even when compared to models at the scale of our largest baseline and Spirit LM (7B), our model exhibits qualitatively different representation-space behavior, consistent with small-scale observations—namely, higher and increasing cross-modal alignment. Based on this and the arguments in Section 1, we expect our contributions to remain beneficial at larger scales.
@@ 7 Limitations
Practical usage of SmolTolk. Since our model is not instruction fine-tuned, its direct applicability to speech LM tasks of interest, such as spoken conversational agents, is limited. Future work could explore instruction tuning and broader evaluations. Additionally, using HuBERT-based tokens limits our model expressivity (Nguyen et al., 2023)—one of speech LMs’ key advantages over cascaded approaches combining text LMs with ASR and TTS. Future work could apply our architecture to model more expressive speech representations.
@@ 8 Conclusions
We argued that early fusion/fission architectures pose limitations for cross-modal transfer in multimodal LLMs. Focusing on text-speech language models, we introduced an alternative architecture based on late fusion/fission and a fission mechanism sensitive to both low- and high-level features. Our SmolTolk models, which embody these ideas, achieved state-of-the-art-level multimodal performance using significantly less data and computational resources than prior work. Representation analyses showed that these components enhance abstraction and improve alignment between speech and text representations. Overall, our results highlight the importance of architectures that respect the compositional hierarchy of multimodal features—an insight that may be key to efficiently adapting text LMs to other perceptual modalities.
@@
Santiago proposed and led the project, defining the hypotheses, model, and experimental design, and developed most of the training, modeling, and analysis codebase. Adel played a key role in data preprocessing and synthesis, text and cross-modal evaluations, and model inference, while also contributing insights on model design. He maintained the codebase with Santiago’s assistance. Ricard provided supervision throughout and contributed significantly to the data pipeline, experimental codebase, and analysis. Yanis assisted with experiment implementation, execution, and discussions. Phil contributed to the current manuscript, discussions, and to Adel’s supervision. Other authors contributed to early-stage discussions and supervision during the JSALT 2024 workshop.
@@ Appendix A Models details
Table 4 describes the SmolLM [Allal et al., 2025] backbones architectural hyperparameters. All SmolTolk models use two layer input and output adapters with the same architecture as the backbone layers. Table 5 describes the resulting models after vocabulary expansion.
@@
As mentioned in Section 3.1, we use a linear layer with bias as the dynamic layer selector $S$. We experimented with simple non-linear MLP selectors; however, these were prone to collapse and resulted in worse overall performance. That said, a more carefully designed non-linear selector could potentially perform better. We also explored alternative ways to define the contextual representation $\mathbf{c}_{i}^{\prime}$ (Equation 3) used as input for the layer selector. Instead of a learned weighted average, we tried concatenating low-dimensional linear projections from each layer’s representations, but this performed more poorly.
@@
For our experiments, we used NVIDIA H100 GPU nodes, each featuring four NVIDIA H100 80GB SXM5 GPUs, dual Intel Sapphire Rapids 48-core processors, 512GB of RAM, and four NVIDIA ConnectX-7 400Gb/s InfiniBand network adapters. SmolTolk-2B used a per-GPU batch size of 8 million tokens and was trained on 64 GPUs across 16 nodes, taking approximately 10 hours to process 32 billion tokens. All models were trained using bfloat16 mixed precision with FlashAttention-2 [Dao, 2024], packed sequences, and PyTorch compile.